# Privacy of the last iterate in cyclically-sampled DP-SGD on nonconvex composite losses

### v1

@@ 1.4 Notation
We consider the swap model for differential privacy. We say two datasets $S$ and $S^{\prime}$ are neighbors, denoted as $S\sim S^{\prime}$, if $S^{\prime}$ can be obtained by swapping one record. A randomized algorithm $\cal A$ is said to be $(\alpha,\varepsilon)$-Rényi differentially private (RDP) if, for every pair of neighboring datasets $S$ and $S^{\prime}$ in the domain of $\cal A$, we have $D_{\alpha}({\cal A}(S)\|{\cal A}(S^{\prime}))\leq\varepsilon$.
2 Randomized proximal Lipschitz operators
This section gives some crucial properties about randomized proximal Lipschitz operators. More specifically, it establishes several RDP bounds based on the closeness of neighboring operators.
We first bound the shifted Rényi divergence of a randomized proximal Lipschitz operator. The proof of this result is a simple extension of the argument in [8, Theorem 22] from $1$-Lipschitz operators to $L$-Lipschitz operators with additive residuals. Nevertheless, we give the full argument in Appendix A.1 for completeness.
Lemma 2.1.
For some $L,\zeta\geq 0$, suppose $\phi^{\prime}$ and $\phi$ satisfy
$\sup_{u}\|\phi^{\prime}(u)-\phi(u)\|\leq s,\quad\|\Phi(x)-\Phi(y)\|\leq L\|x-y\|+\zeta,\quad\forall\Phi\in\{\phi,\phi^{\prime}\},$ (4)
for any $x,y\in\operatorname*{dom}\phi\cap\operatorname*{dom}\phi^{\prime}$. Moreover, let $Z\sim{\cal N}(0,\sigma^{2}I)$ and $\psi\in\overline{{\rm Conv}}\,(\mathbb{R}^{n})$. Then, for any scalars $a,\tau\geq 0$ and $\alpha\in(1,\infty)$ satisfying $L\tau+\zeta+s-a\geq 0,$ and random variables $Y$ and $Y^{\prime}$, it holds that
Note that the second inequality in (4) is equivalent to $\Phi$ being $L$-Lipschitz continuous when $\zeta=0$, and the conditions in (4) need to only hold on $\operatorname*{dom}\phi\cap\operatorname*{dom}\phi^{\prime}$.
We now apply (5) to a sequence of points generated by the randomized proximal update $Y\xleftarrow{}\operatorname{prox}_{\psi}(\phi(Y)+Z)$, under different assumptions on $\zeta$ and $\tau$. This particular result generalizes the one in [8], which only considers the case of $L=1$ and $\zeta=0$. For conciseness, we leave the full proof in Appendix A.2.
Before proceeding, we define the following useful quantities:
$\theta_{L}(0)=0,\quad\theta_{L}(s):=\frac{L^{2(s-1)}}{\sum_{j=0}^{s-1}L^{2j}}\quad\forall s\geq 1.$ (6)
Proposition 2.2.
Let $L,\zeta\geq 0$, $T\geq 1$, and $\ell\in[T]$ be fixed. Given $\psi\in\overline{{\rm Conv}}\,(\mathbb{R}^{n})$, suppose $\{\phi_{t}\}$, $\{\phi_{t}^{\prime}\}$, and $\bar{s}>0$ satisfy (4) with
$\begin{gathered}Y_{t}:=\operatorname{prox}_{\psi}(\phi_{t}(Y_{t-1})+Z_{t}),\quad Y_{t}^{\prime}:=\operatorname{prox}_{\psi}(\phi_{t}^{\prime}(Y_{t-1}^{\prime})+Z_{t}),\quad\forall t\geq 1\end{gathered}$
Then, for any $\tau\geq 0$ and $E=\lfloor T/\ell\rfloor$, the following statements hold:
  • (a)
    if $\zeta=0$, then
    $D_{\alpha}(Y_{T}\|Y_{T}^{\prime})-D_{\alpha}^{(\tau)}(Y_{0}\|Y_{0}^{\prime})$
  • (b)
    if $\tau=0$ and $\zeta=\bar{s}$, then
Some remarks about Proposition 2.2 are in order. First, part (a) shows that if $\phi_{t}$ and $\phi_{t}^{\prime}$ only differ at $t=1$, then $D_{\alpha}(Y_{T}\|Y_{T}^{\prime})$ is finite for any $T$. Second, part (a) also shows that if $\phi_{t}$ and $\phi_{t}^{\prime}$ differ cyclically for a cycle length of $\ell$, then the divergence between $Y_{T}$ and $Y_{T}^{\prime}$ grows linearly with the number of cycles $E$. Third, part (b) gives a bound that is independent of $L$. Finally, both of the bounds in parts (a) and (b) can be viewed as Rényi divergences between the Gaussian random variables ${\cal N}(0,\sigma^{2}I)$ and ${\cal N}(\mu,\sigma^{2}I)$ for different values of $\mu$.
In Appendix B, we give a detailed discussion of how the residuals $a$ from Lemma 2.1 are chosen to prove Proposition 2.2(a). In particular, we prove that the chosen residuals yield the tightest RDP bound that can achieved by repeatedly applying (5).
3 Curvature assumptions and operator analysis
This section states some curvature assumptions for future reference and develops the properties of an SGD-like update.

### v3

@@ 1.1 Background
We consider the swap model for differential privacy. We say two datasets $S$ and $S^{\prime}$ are neighbors, denoted as $S\sim S^{\prime}$, if $S^{\prime}$ can be obtained by swapping one record. A randomized algorithm $\cal A$ is said to be $(\alpha,\varepsilon)$-RDP if, for every pair of neighboring datasets $S$ and $S^{\prime}$ in the domain of $\cal A$, we have $D_{\alpha}({\cal A}(S)\|{\cal A}(S^{\prime}))\leq\varepsilon$.
$\cal A$} satisfies local DP if for all records $x_{i}$ and $x_{j}$, $D_{\alpha}({\cal A}(x_{i})\|{\cal A}(x_{j}))\leq\varepsilon$. Finally, we use the following variable conventions: $\ell$ is the number of batches (or iterations) in a dataset pass, $E$ is the number of dataset passes, $T=E\cdot\ell$ is the total number of iterations, $k$ is the total number of per-example losses, $b$ is the batch size, $\lambda$ is the DP-SGD stepsize, and $C$ is the clipping norm.
DP-SGD variant. Algorithm 1 outlines the specific variant of DP-SGD applied to (1). This variant takes as input $k$ per-example loss functions $\{f_{i}\}^{k}_{i=1}$, the number of iterations $T$, iid samples $\{N_{t}\}^{T}_{t=1}$ from a spherical Gaussian distribution ${\cal N}(0,\sigma^{2}I)$, initial model weights $X_{0}$, batch size, stepsize, and $\ell_{2}$-clipping norm values. Model weights are updated as follows. At time step $t$, the algorithm (i) selects a batch of examples by cyclically traversing $\{f_{i}\}^{k}_{i=1}$, (ii) computes the average $g_{t}$ of clipped per example gradients at $X_{t-1}$, and (iii) updates $X_{t-1}$ using a noisy gradient. Finally, the algorithm returns the last iterate, $X_{T}$.
Input: $\{f_{i}\}_{i=1}^{k}$, $h$, iid samples $\{N_{t}\}\subseteq\mathbb{R}^{n}$ from ${\cal N}(0,\sigma^{2}I)$, $X_{0}\in\operatorname*{dom}h$;
Data: batch size $b$, stepsize $\lambda$, clipping norm $C$, iteration limit $T$, steps per dataset pass $\ell$;
Output: $X_{T}\in\operatorname*{dom}h$;
0.3em for $t=1,\ldots,T$ do
 $j_{t}\leftarrow b(t-1)\bmod k$
 $B_{t}\leftarrow\{j_{t}+1,\ldots,j_{t}+b\}$
 $g_{t}\leftarrow(1/b)\sum_{i\in B_{t}}{\rm Clip}_{C}(\nabla f_{i}(X_{t-1}))$
 $X_{t}\leftarrow\operatorname{prox}_{\lambda h}\left(X_{t-1}-\lambda g_{t}+N_{t}\right)$
end for
return $X_{T}$
Algorithm 1 Cyclically-traversed last-iterate DP-SGD
1.2 Outline of approach
We now outline our approach of tackling the problem of interest. A more formal treatment is given in Section 2.
To motivate our approach, we provide a brief overview of previous well-known methods. An early approach of analyzing Algorithm 1 is to develop a bound based on local DP for a single dataset pass and extend this bound for multiple passes (see, for example, [27, 14]). While straightforward, this approach can be overly restrictive in a centralized setting. Privacy Amplification by Subsampling (PABS) (see subsequent work [13] for a comparison of different sampling methods) improves on the previous approach in certain regimes. Although this method allows for clean privacy accounting, its reliance on Poisson subsampling makes it impractical for large-scale applications (see Appendix D for an extended discussion). The work started by [17] addressed the limitations of PABS with Privacy Amplification by Iteration (PABI), which achieves a bound for releasing the final DP-SGD iterate. This PABI bound improves the baseline bound in certain regimes incorporating a contraction factor dependent on the loss function’s convexity parameters.
Our approach is inspired by PABI, but relaxes several of its convexity assumptions. For added context, we briefly review PABI below. Under the assumption that the loss function of (1) is convex and $Q$-Lipschitz, and that $h$ is the indicator of a closed convex set, [17] shows that the DP-SGD update in Algorithm 1 with small constant stepsize $\lambda$ and no gradient clipping is a nonexpansive operator. This property can be then combined with the following technical result about nonexpansive operators.
Theorem 1.1.
Suppose we are given iterates $\{X_{t}\}$ and $\{X_{t}^{\prime}\}$, nonexpansive operators $\{\phi_{t}\}$ and $\{\phi_{t}^{\prime}\}$, iid Gaussian random variables $\{N_{t}\}$, and scalars $\{s_{t}\}$ satisfying
$X_{t}=\phi_{t}(X_{t-1})+N_{t},\quad X_{t}^{\prime}=\phi_{t}^{\prime}(X_{t-1}^{\prime})+N_{t},\quad\sup_{x}\|\phi_{t}(x)-\phi_{t}^{\prime}(x^{\prime})\|\leq s_{t}\quad\forall t\in[T].$
For any scalar sequences $\{a_{t}\}$ and $\{z_{t}\}$ satisfying
$z_{t}=\sum_{i\leq t}s_{i}-\sum_{i\leq t}a_{i}\geq 0,\quad z_{t}\geq 0,\quad a_{t}\geq 0,\quad\forall t\in[T],$ (6)
we obtain the following last-iterate shifted Rényi divergence bound:
$D_{\alpha}^{(z_{T})}(X_{T}\|X_{T}^{\prime})-D_{\alpha}(X_{0}\|X_{0}^{\prime})\leq\frac{\alpha}{2\sigma^{2}}\sum_{t=1}^{T}a_{t}^{2}=:{\cal R}_{T}(\{a_{t}\})\quad\forall T\geq 1,\quad\forall\alpha\geq 1.$ (7)
More specifically, assuming that the DP-SGD iterates first differ at index $t^{*}$, i.e., $X_{t^{*}}^{\prime}\neq X_{t^{*}}$ and $X_{t}^{\prime}=X_{t}$ for every $t<t^{*}$, the operators $\{\phi_{t}\}$ and $\{\phi_{t}^{\prime}\}$ in the above theorem can be formed with $s_{t^{*}}=2\lambda Q$ and $s_{t}=0$ for every $t\neq t^{*}$. Consequently, one can select $\{a_{t}\}$ so that the shift satisfies $z_{T}=0$ and obtain a closed form bound ${\cal R}_{T}(\{a_{t}\})=\Theta(\alpha[\lambda Q/\sigma]^{2})$ in (7), which yields a corresponding RDP bound when $X_{0}=X_{0}^{\prime}$. The generalization to multiple dataset passes follows similarly but the final bound scales with the number of dataset passes $E$. Specifically, we have ${\cal R}_{T}(\{a_{t}\})=\Theta(\alpha[E/\ell]\cdot[\lambda Q/\sigma]^{2})$.
In the more practical setting where (i) the loss function in (1) is nonconvex and not necessarily $Q$-Lipschitz, (ii) gradient clipping is applied, and (iii) $h$ is nonsmooth, it is no longer clear to what extent the corresponding DP-SGD operators $\{\phi_{t}\}$ and $\{\phi_{t}^{\prime}\}$ are nonexpansive or how $\{s_{t}\}$ should be obtained. Furthermore, the first inequality of (6) no longer holds, and additional technical issues arise when analyzing the case of multiple dataset passes.
Our approach. We generalize the above argument and combine it with additional analyses of weakly-convex functions and proximal operators to relax several strong assumptions. A sketch of our approach is given below, and formal arguments can be found in subsequent sections.
General operator analysis. In Lemma A.2, we study properties of operators $\phi$ and $\phi^{\prime}$ satisfying
$\sup_{u}\|\phi^{\prime}(u)-\phi(u)\|\leq s,\quad\|\Phi(x)-\Phi(y)\|\leq L\|x-y\|+\zeta,\quad\forall\Phi\in\{\phi,\phi^{\prime}\}.$ (8)
Note that if $\Phi$ is Hölder continuous, then it can be shown [22] that it satisfies the second inequality in (8) [^7] More specifically, if $\Phi$ is $\eta$-Hölder continuous with modulus $H$, then (8) holds with $L=H\rho/2$ and $\zeta=H\eta([1-\eta]/\rho)^{(1-\eta)/\eta}$ for any $\rho>0$.. Using these properties, we then establish in Proposition A.5 that if $\{Y_{t}\}$ and $\{Y_{t}^{\prime}\}$ are generated by a specific sequence of randomized proximal operators using $\phi$ and $\phi^{\prime}$, respectively, then (roughly)
$D_{\alpha}(Y_{T}\|Y_{T}^{\prime})-D_{\alpha}(Y_{0}\|Y_{0}^{\prime})\preceq\frac{\alpha}{\sigma^{2}}\cdot{\cal{F}}(T),$
[…]
Lipschitz properties of the DP-SGD update. Denoting ${\cal A}_{\lambda}(\cdot)$ as the DP-SGD update function [^8] Or any SGD-like update as in (12)., i.e., $X_{t}={\cal A}_{\lambda}(X_{t-1})$ for every $t$ in Algorithm 1, we show in Proposition 2.1 that — depending on our assumptions on $h$ and stepsize $\lambda$ — the operator ${\cal A}_{\lambda}(\cdot)$ satisfies the second inequality of (8) with $\Phi=A_{\lambda}$ for different values of $L$ and $\zeta$.
More specifically, when the domain of $h$ is bounded, we have $(L,\zeta)=(1,2\lambda C)$ in (8) for clipping norm $C$. On the other hand, when $\lambda$ is sufficiently small, we have $\zeta=0$ and $L$ being a constant that (i) tends to $\sqrt{2}$ when the weak convexity parameter $m$ tends to zero, i.e., $f$ becomes more convex; and (ii) tends to one when no clipping is performed and $m$ in (i) tends to zero. This continuity with respect to the weak convexity parameter appears to be new, and it is proved in Appendix A.2 by using topological properties about weakly-convex functions and proximal operators.
Privacy bounds for DP-SGD. For neighboring DP-SGD iterates $X_{T}$ and $X_{T}^{\prime}$ , we combine the above results in Theorems 2.2 and 2.3 to obtain RDP bounds of the form
$D_{\alpha}(X_{T}\|X_{T}^{\prime})\preceq\frac{\alpha}{\sigma^{2}}\cdot{\cal{B}}_{\lambda}(C,b,T,\ell),$ (9)
where $C$, $b$, $T$, $\ell$, are as in Algorithm 1. More specifically, assuming that the DP-SGD iterates are contained within an $\ell_{2}$ ball of diameter $d_{h}$ and each $\nabla f_{i}$ is Lipschitz continuous, we obtain (9) with $B_{\lambda}(C,b,T,\ell)=(L_{\lambda}d_{h}+\lambda C/b)^{2}$ for for some $L_{\lambda}=\sqrt{1+\kappa\lambda m}$, $\kappa\leq 4$, and small enough $\lambda$. When the iterates are (possibly) unbounded, we obtain (9) with
$B_{\lambda}(C,b,T,\ell)=\left(1+\left[\frac{T}{\ell}\right]\left[\frac{L^{2\ell}_{\lambda}}{\sum_{i=1}^{\ell}L^{2i}_{\lambda}}\right]\right)\left(\frac{\lambda C}{b}\right)^{2}.$ (10)
1.3 Related work
We first present high-level descriptions of related works in the convex and nonconvex settings, followed by more general works that use advanced composition to obtain loose bounds on the last iterate. We then conclude with some summary tables and figures, and a discussion of technical nuances that carefully compares our work to existing literature.
Convex setting. Given the challenge of proving tight bounds in the general setting, a number of prior analyses focus on the convex case. Works by [17, 16] additionally assume Lispchitz continuity of the loss function to obtain results. The work of [12] studies multiple passes over the data, but their results only apply to the smooth, strongly convex, and full batch setting without clipping. The work of [28] improves these results and extend them to mini-batches both with “shuffle and partition” and “sampling without replacement” strategies. Similarly, results in by [2], and its extension [3], consider only convex Lipschitz smooth functions. The contemporary work of [8] introduces the shifted interpolated process under $f$-DP, allowing for tight characterizations of privacy by iteration for Rényi and other generalized DP definitions.
Notice that none of the above works study clipping, all assume access to the Lipschitz constant of the loss function, and require convexity, limiting their practical viability.
Nonconvex setting. We now discuss papers that do not require convexity of the loss function. [4] analyze the privacy loss dynamics for nonconvex functions, but their analysis differs from ours in two ways. First, they assume that their DP-SGD batches are obtained by Poisson sampling or sampling without replacement. Second, their results require numerically solving a minimization problem that can be hard in practice.
A contemporary work [^9] Appearing after the first version of this preprint. by [11] derives bounds under the assumption that the loss gradient is Hölder continuous and the loss function is Lispchitz continuous when it is also convex. However, this work needs an additional assumption, that constants $L>0$ and $\eta\in(0,1]$ satisfying $\|\nabla f(x)-\nabla f(y)\|\leq L\|x-y\|^{\eta}$ are known and used in a specific optimization subproblem. While [11] focuses on tight theoretical bounds under specific conditions (such as the full batch and single-epoch setting), we prioritize bridging the gap between theory and practice by addressing the complexities of real-world deployments.
Non-specialized analyses. Prior works on DP-SGD accounting often rely on loose bounds that allow for the release of intermediate updates [20, 6, 1]. These works rely on differential privacy advanced composition results [19], resulting in noise with standard deviation that scales as $\sqrt{T}$ [6, 1]. Alternatively, using disjoint batches decreases the dependence from the number of iterations $T$ to the number of epochs $E$ (see the first row in Table 2). However, the assumption that all intermediate updates are released can be stringent for certain loss regimes [17, 16], where this bound can be contracted based on loss smoothness and convexity parameters, if only the last iterate is released.
While our work focuses on the privacy guarantees of DP-SGD, it’s important to acknowledge the parallel research efforts exploring the convergence properties of shuffling methods. Recent studies, such as [23] and [9], have established convergence bounds for strongly convex and/or smooth functions in various settings, albeit without considering privacy. These works provide valuable insights into the optimization behavior of shuffling techniques, which can inform future research at the intersection of privacy and optimization.
Summary tables and graphs. Table 1 lists and labels some assumptions that are commonly used in the RDP literature. Table 2 compares our bounds against other RDP bounds. Note that the multi-epoch noisy-SGD algorithm in [17, Algorithm 1] only considers the case where the number of dataset passes equals the number of batches in dataset pass and does not consider batched gradients. As a consequence of the latter, its corresponding RDP upper bound does not depend on the batch size $b$. Figure 1 compares our bounds against other bounds as function of the number of iterations performed, under various settings. Note that we do not consider the multi-epoch bound in [11, Theorem A.1] as it requires solving $T$ nonlinear programming problems, and we do not compare the with the bound in [25] because it exploits the fact that batches are randomly sampled (whereas our bounds assume the input batches must be obtained deterministically).
Label Description
${\cal I}_{c}$ The regularizer $h$ is the indicator of a closed convex set
[…]
${\cal B}_{1}$ input data batches are of size one
$\cal D$} input data batches are disjoint
$\cal P$} input data batches are obtained using Poisson subsampling with sampling rate $1/\ell$
${\cal S}_{Q}^{0}$ $f_{i}$ is $Q$-Lipschitz continuous for every $i\in[k]$, and $Q$ is known
${\cal S}_{L}^{1}$ $\nabla f_{i}$ is $L$-Lipschitz continuous for every $i\in[k]$, and $L$ is known
${\cal R}_{H,\eta}^{1}$ $\nabla f_{i}$ is $(H,\eta)$-Hölder continuous for every $i\in[k]$, and $H$ and $\eta$ are known
$\Lambda$ the stepsize needs to be small relative to certain smoothness constants
$\cal A$} the given RDP bounds only hold for a small range of values of $\alpha$ and $\sigma$
$\cal N$} no gradient clipping is applied or the global sensitivity is known
$\tilde{{\cal N}}$ when $\phi$ is convex, no gradient clipping is applied or the global sensitivity is known
Table 1: List of common assumptions used in RDP accounting literature. For conciseness we let $\ell$ denote the number of iterations/batches in a dataset pass and $\operatorname*{dom}h$ denote the domain of $h$.
Source Asymptotic $(\alpha,\epsilon)$-RDP upper bounds Assumptions
Convex $\phi$ Nonconvex $\phi$ (see Table 1)
[14, Theorem 3.1] [^10] The bound follows directly from Theorem 8 in [5] $\dfrac{\alpha E}{2}\left[\dfrac{\lambda C}{\sigma b}\right]^{2}$ same as convex $\cal D$}
[…]
[17, Theorem 35] $\dfrac{\alpha E}{\ell}\left[\dfrac{\lambda Q}{\sigma}\right]^{2}$ none ${\cal I}_{c},{\cal N},\Lambda,S_{Q}^{0},S_{L}^{1},{\color[rgb]{1,0,0}{\cal B}_{1}}$
[…]
[11, Theorem A.1] numerical procedure [^12] Obtained by solving $T$ nonlinear programming problems. numerical procedure [^13] Same procedure as in the nonconvex case, but with different parameters. ${\cal I}_{c},\tilde{{\cal N}},\Lambda,{S}_{Q}^{0},{\cal R}_{H,\eta}^{1},{\cal H}$
Ours $\dfrac{\alpha}{\sigma^{2}}\left[L_{\lambda}d_{h}+\dfrac{\lambda C}{b}\right]^{2}$ same as convex $\Lambda,S_{L}^{1},{\cal H},{\cal D}$
$\dfrac{\alpha E}{\ell}\left[\dfrac{\lambda C}{b\sigma}\right]^{2}$ $\alpha E\theta_{L_{\lambda}}(\ell)\left[\dfrac{\lambda C}{\sigma}\right]^{2}$ $\Lambda,S_{L}^{1},{\cal N},{\cal D}$
$\alpha E\theta_{\sqrt{2}}(\ell)\left[\dfrac{\lambda C}{\sigma}\right]^{2}$ $\alpha E\theta_{\sqrt{2}L_{\lambda}}(\ell)\left[\dfrac{\lambda C}{\sigma}\right]^{2}$ $\Lambda,S_{L}^{1},,{\cal D}$
Table 2: Asymptotic $\varepsilon$ upper bounds for $(\alpha,\varepsilon)$-Rényi differential privacy of the last iterate after $T$ iterations of DP-SGD. Here, $\lambda$ is the stepsize, $C$ is the clipping norm, $\sigma$ is the standard deviation of the Gaussian noise, $\ell$ is the number of iterations in one dataset pass, and $E$ is the number of dataset passes. Also, $L_{\lambda}:=\sqrt{1+\lambda\kappa m}$ for some $\kappa\leq 4$ and weak-convexity parameter $m$ (see (11)), and $\theta_{L}(\ell):=L^{2(\ell-1)}/\sum_{i=0}^{\ell-1}L^{2i}$. Particularly strong assumptions are highlighted in red.
Figure 1: Log-scale last-iterate RDP bounds for Algorithm 1 as a function of number of iterations. The fixed algorithmic parameters are $\lambda=10^{-5}$, $C=10$, $\sigma=10^{-5}$, $T=10^{5}$, $b=10^{1}$, and $k=10^{4}$. The free parameters $m$, $M$, and $Q$ are the weak convexity, Lipschitz-smoothness, and Lipschitz constants of $f$, respectively, while $C$ is the clipping norm. The left plot considers weakly-convex losses under the settings of no gradient clipping. The right plot considers the convex case where $Q$ may differ from $C$. The RDP composition bounds are from [20], the PABI bounds are from [17], and the local DP bounds are from [14].
Technical nuances. The bound in (9) with ${\cal B}_{\lambda}$ as in (10) and $L_{\lambda}=1$ might appear to follow from subsampled RDP composition results such as [25]. However, those results only apply to DP-SGD variants where the batches are sampled randomly, an assumption that does not hold when batches are cyclically traversed. While established Python libraries like Opacus [29] and TensorFlow Privacy [24] implement and account for random sampled batches (such as those obtained by Poisson subsampling), these implementations address a different issue. One has to ensure the optimizer truly samples at random from a pre-specified distribution, which becomes incredibly difficult with large-scale datasets (see Appendix D for an extended discussion). Consequently, any privacy guarantee predicated on this idealized random sampling assumption becomes effectively meaningless when the actual optimization process deviates from it (as is the case with cyclical batch traversal). It is worth mentioning that DP-FTRL [18] was specifically developed to address this gap and the method accepts a potentially weaker DP guarantee in exchange for practical applicability.
For the special case where (i) only one dataset pass is performed, (ii) the objective function is nonconvex , and (iii) each $\nabla f_{i}$ is Lipschitz continuous, the bound in (9) with ${\cal B}_{\lambda}$ as in (10) holds with $T=\ell$. Consequently, we obtain an RDP bound that does not grow with the number of iterations — in contrast to the RDP composition bounds in [20] which do scale linearly in $T$ or $E$, depending on the sampling assumption. Note that (16) and Theorem 2.3 show that as $\theta_{L_{\lambda}}(\cdot)\downarrow 1$, our bounds converge to an expression that depends linearly on $E/\ell$, matching the bound for PABI. Figure 1 illustrates the regimes under which we improve previous work.
The addition of the composite term $h(\cdot)$ in (1) is not a trivial extension and greatly complicates the analysis. For example, the objective function $\phi$ in (1) is no longer differentiable in general (even on $\operatorname*{dom}\phi$), and more general analyses must be used to handle this nonsmoothness. Under stepsize $\lambda$ and $L$-Lipschitz continuous $\nabla f_{i}$ for $i\in[k]$, paper [11] shows that the DP-SGD update in the nonconvex case is $(1+\lambda L)$-Lipschitz continuous, which is independent of the weak convexity constant. Consequently, the established RDP bounds in [11] in the setting where the weak convexity constant $m$ is positive but near zero (i.e., the function is nearly convex) may be an overestimate of the true RDP bound.
For the convex case, it is worth emphasizing that we do not require each to be Lipschitz continuous case in order to bound $\nabla f_{i}$ (see, for example, [17, 2, 16, 11] which do require this assumption). As a consequence, our analysis is applicable to a substantially wider class of objective functions. Moreover, all other existing bounds in the literature of the form in (9) replace the parameter $C$ with a Lipschitz constant $Q$ of $\phi(x)$ from (1), and these bounds are generally proportional to $Q^{2}$. Consequently, when $C\ll Q$, e.g., when $\phi(x)$ is a quadratic function on a large compact domain, our bounds are significantly tighter (see Figure 1 for an illustration).
Organization. The remainder of the paper gives a formal presentation of the results, including the key assumptions on (1), the topological properties of the DP-SGD update operator, and the non-asymptotic RDP bounds on the last DP-SGD iterates.
2 Privacy bounds for DP-SGD
This section formally presents the main RDP bounds for the last iterates of Algorithm 1. For conciseness, the lengthy proofs of the main results are given in Appendix A.